OpenStudy (anonymous):

simplify √3w^(2)x . √6w^(2)x

OpenStudy (mertsj):

\[\sqrt{3w^2x}\sqrt{6w^2x}=\sqrt{18w^4x^2}=\sqrt{9(2)w^4x^2}=3w^2x \sqrt{2}\]